The half-life of a certain radioactive material is 42 days. An initial amount of the material has a mass of 49 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 8 days. Round your answer to the nearest thousandth.
Answer by ankor@dixie-net.com(22740)

The half-life of a certain radioactive material is 42 days.
An initial amount of the material has a mass of 49 kg.
Write an exponential function that models the decay of this material.
:
A = Ao*2^(-t/h), where:
A = amt after t time
Ao = initial amt (t=0)
t = time of decay
h = half life of substance
:
A = 49*2^(-t/42)
:
:
Find how much radioactive material remains after 8 days.
A = 49*2^(-8/42)
A = 49*.8763
A = 42.9395 ~ 42.940, to the nearest thousandth.